Debra is six years older than Raymond if the product of their ages is 55 how old is Debra
Answer by ankor@dixie-net.com(22740)

Debra is six years older than Raymond
d = r + 6
if the product of their ages is 55
d * r = 55
replace d with (r+6)
(r+6)*r = 55
a quadratic equation
r^2 + 6r - 55 = 0
Factors to
(r+11)(r-5) = 0
positive solution is what we want here
r = 5 yrs old is Raymond
then "how old is Debra"
d = 5 + 6 = 11 yrs